Kirtland Corporation uses a periodic inventory system. At the end of the annual accounting period, December 31, the accounting records for the most popular item in inventory showed the following:

Transactions Units Unit Cost
Beginning inventory, January 1 390 $6.00
Transactions during the year:
a. Purchase, January 30 290 3.30
b. Purchase, May 1 450 7.00
c. Sale ($8 each) (150)
d. Sale ($8 each) (690)


Required:

a. Compute the amount of goods available for sale.

b. & c. Compute the amount of ending inventory and cost of goods sold at December 31, under Average cost, First-in, first-out, Last-in, first-out and Specific identification inventory costing methods. For Specific identification, assume that the first sale was selected two-fifths from the beginning inventory and three-fifths from the purchase of January 30. Assume that the second sale was selected from the remainder of the beginning inventory, with the balance from the purchase of May 1.

Complete this question by entering your answers in the tabs below.

  • Req A
  • Req B and C

Compute the amount of ending inventory and cost of goods sold at December 31 under Average cost, First-in, first-out, Last-in, first-out, Specific identification of the inventory costing methods. For Specific identification, assume that the first sale was selected two-fifths from the beginning inventory and three-fifths from the purchase of January 30. Assume that the second sale was selected from the remainder of the beginning inventory, with the balance from the purchase of May 1. (Do not round intermediate calculations. Round your answers to the nearest whole dollar amount.)

Show less

Average Cost First-In, First-Out Last-In, First-Out Specific Identification
Ending inventory $1,656
Cost of goods sold $4,796
Answer #1
a] Transaction Units Unit Cost Total Cost
January 1, Beginning inventory 390 $                  6.00 $             2,340
January 30, Purchase 290 $                  3.30 $ 957
May 1, Purchase 450 $                  7.00 $             3,150
Total 1130 $             6,447
Cost of goods available for sale = $             6,447
Sale in units = 150+690 = 840
Ending inventory in units 290
Average unit cost = 6447/1130 = $                5.71
b & c] Average cost:
Ending inventory = 290*5.71 = $             1,655
COGS = 6447-1655 = $             4,792
FIFO:
Ending inventory = 290*7 = $             2,030
COGS = 6447-2030 = $             4,417
LIFO:
Ending inventory = 290*6 = $             1,740
COGS = 6447-1740 = $             4,707
Specific identification:
Ending inventory = 200*3.3+90*7 = $             1,290
COGS = (150*6*2/5+150*3.3*3/5)+(330*6+360*7) = $             5,157